Session: Feb 1999 (session 1)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Zacchiroli at Sat Feb 13 02:40:23 MET 1999) Success
No problem during the observations

Onsala (A. Polatidis at Sat Feb 13 09:22:32 MET 1999) Minor failures
Due to power failure at the station lost data between 0020-0238 UT. During the last scan the telescope was unwraping the cable in azimuth, hence no valid data.

Westerbork (Tony Foley at Sun Feb 14 16:23:15 MET 1999) Minor failures
No known problems. We used a single dish setup (RT7)

Jodrell (Alastair Gunn at Sun Feb 14 18:15:21 MET 1999) Success
No known problems.

Torun (K.Borkowski at Mon Feb 15 13:09:44 MET 1999) Success
No known problems.

Noto (C.Trigilio at Tue Feb 16 18:15:05 MET 1999) Success

Effelsberg (dgraham at Mon Feb 22 15:18:06 MET 1999) Success

Calibration (Jean\-Francois Desmurs at Wed Mar 3 12:05:04 MET 1999) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the feb99 directory under the name: v129c4.antab. Read the comments in the file to get further information.